OpenStudy (anonymous):

Give exact and approximate solutions to three decimals place. x^2+18x+81=1

jhonyy9 (jhonyy9):

- this is (x+9)*2 =1 --- x+9=1 --- x=-8 64-144+80=0

OpenStudy (anonymous):

ok so what are the decimals I am confused?

jhonyy9 (jhonyy9):

have not decimals

OpenStudy (anonymous):

it is asking for exact answer then in form i ?

OpenStudy (anonymous):

so i put 0

jhonyy9 (jhonyy9):

here you not will get i never - complex numbers like solution

jhonyy9 (jhonyy9):

but now i see that is again right solution -10 too

jhonyy9 (jhonyy9):

because -10 +9=(-1)*2 =1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this one is the same (x+7)^2=64

jhonyy9 (jhonyy9):

so in finaly we can seeing that there are two solutions -10 and -8

OpenStudy (anonymous):

r we on the first or second one

jhonyy9 (jhonyy9):

here in the secongd equation you have (x+7)*2 = 64 = 8*2 - so if you make radical from both sides you get x+7=8 --- x=1

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so x=1

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

hope so much that i have helped to you and that all is clearly

OpenStudy (anonymous):

yes I have some more you want to help me with those too please

jhonyy9 (jhonyy9):

- for you being sure rigth in solutions you need to make probe

jhonyy9 (jhonyy9):

ok let's go

OpenStudy (anonymous):

Determine the nature of the solution if the equation y^2= 2/3 y+ 3/7 the y is with the whole 2/3 not just the 3

OpenStudy (anonymous):

one real solution? two non- real solution? two real solution?

OpenStudy (anonymous):

multiply choice here

jhonyy9 (jhonyy9):

please us parentheses

OpenStudy (anonymous):

y^2=(2/3y)+ (3/7)

jhonyy9 (jhonyy9):

y*2 = 2/3y +3/7 what is the denominator common ?

OpenStudy (anonymous):

the y is with the whole (2/3) not just the 3

OpenStudy (anonymous):

21

jhonyy9 (jhonyy9):

i want to make you understand clearly

jhonyy9 (jhonyy9):

why 21 ?

jhonyy9 (jhonyy9):

what are there like denominators ?

OpenStudy (anonymous):

7*3=21 3*7=21 that is the closest they have togather

jhonyy9 (jhonyy9):

and y there not is ?

jhonyy9 (jhonyy9):

becaus eyou have wrote 2/3y and not 2y/3

OpenStudy (anonymous):

I do not understand what you are asking?

OpenStudy (anonymous):

y is with the whole fraction 2/3 I am not sure how to write other then 2 ----y 3

jhonyy9 (jhonyy9):

how i have wrote to you before this you can write 2y/3

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

this is clearly ?

OpenStudy (anonymous):

yes!

jhonyy9 (jhonyy9):

if yes what is the denominator common ?

jhonyy9 (jhonyy9):

ok 21 and in this case you need multiply the first term with ?

OpenStudy (anonymous):

I still am getting 21! Is that not correct or should it be 21y

jhonyy9 (jhonyy9):

ok yes is correct than y not is in denominator is 21

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

and now how you will continue ?

OpenStudy (anonymous):

so 21 is the denominator?

jhonyy9 (jhonyy9):

if the denominator common is 21 what you need to do ?

jhonyy9 (jhonyy9):

multiplie the first term with ?

jhonyy9 (jhonyy9):

if you have y*2 =2y/3 +3/7

jhonyy9 (jhonyy9):

because y*2 have denominator 1 you need to multiplie this term with 21

jhonyy9 (jhonyy9):

this is clearly ?